A fly wheel of inertia 0.32kgm^2 is rotated steadily at 120rad/s by a 50watt electric motor. Calculate the kinetic energy of the fly wheel and the angular momentum of the fly wheel.
1
Expert's answer
2020-10-19T13:25:36-0400

"E=\frac{I\omega^2}{2}=\frac{0.32\cdot120^2}{2}=2304" J.


"L=pR=mvR=m\omega R^2=2\cdot\frac{1}{2}mR^2\omega=2I\omega=2\cdot0.32\cdot120=76.8 \frac{kg\cdot m^2}{s}"